A motorist with an expired license tag travels at a constant speed of 14.4 m/s down a street, and a policeman on a motorcycle, taking another 4.01 s to finish his donut, gives chase at an acceleration of 2.23 m/s2.

(a) Find the time required to catch the car.
t = s

(b) Find the distance the trooper travels while overtaking the motorist.
x = m
Answer #1

Solution)

Relative distance between two when policemen starts = 14.4 * 4.01= 57.74 m

Given, relative acceleration = 2.23 m/s^2

initial relative velocity = -14.4 m/s

Now, we know

s = ut + at/2^2

507.74= -14.4t + 2.23t/2^2

1.115 t^2 -14.4t- 50.74= 0

t = 15.95sec (Ans)

======

b)s = ut + at^2/2

u = 0

a = 2m/s2

t = 15.95 sec

s = 254.40 m(Ans)

======

Good luck!:)